OpenStudy (anonymous):

The surface area of a sphere is 196 pi inches squared. How do you find the volume of the hemisphere?

OpenStudy (isaidavila):

Surface área of a sphere is SA=4πr^2 so 196π=4πr^2

OpenStudy (jdoe0001):

\(\bf \textit{surface area of a sphere}={\color{brown}{ 4\pi r^2 }}\qquad {\color{brown}{ surface }} =196\pi \\ \quad \\ thus\qquad 4\pi {\color{blue}{ r}}^2=196\pi \implies {\color{blue}{ r}}=? \\ \quad \\ \textit{keep in mind that }volume = \cfrac{4}{3}\pi {\color{blue}{ r}}^3\)

OpenStudy (jdoe0001):

so.. find "r" and plug it in the volume formula

OpenStudy (anonymous):

r = 6.96419413859?

OpenStudy (anonymous):

1414.82008731 divided by 2 is the answer???

OpenStudy (mathmale):

Would someone take a moment and explain what a "hemisphere" is? That will affect the outcome of this problem.

OpenStudy (jdoe0001):

hhhmm ohh ... a Hemi's... hmm thought it was the whole thing =)

OpenStudy (anonymous):

I know it is half of a sphere, but would dividing the volume by 2 find the volume of the hemisphere?

OpenStudy (jdoe0001):

either way, you'll need to find "r"

OpenStudy (jdoe0001):

tsunekawa yes

OpenStudy (mathmale):

Yes: a hemisphere is simply half a sphere. Use this fact as you come up with a solution to this posted problem.

OpenStudy (anonymous):

Therefore, the answer is about 707.4?

OpenStudy (anonymous):

Oh, thank you

OpenStudy (mathmale):

My pleasure!

OpenStudy (anonymous):

Wait, but 707.4 is correct?